LemonDragon101:

Given cos(α) = -3/14, with α in quadrant III, and sin(β) = -12/23, with β in quadrant IV. Find tan(α + β)

Uniful3:

(√71295 - 36) / (3 + 12√187)

Uniful3:

is it correct..?

Aratox:

This is a long answer explaining everything that you may need, your desired/final answer is at the bottom. We can begin by finding the value of sin(α) using the Pythagorean identity sin^2(α) + cos^2(α) = 1. Since cos(α) = -3/14, we have (-3/14)^2 + sin^2(α) = 1. Simplifying, we get 9/196 + sin^2(α) = 1. Subtracting 9/196 from both sides we find sin^2(α) = 187/196. Taking the square root of both sides and remembering that α is in quadrant III (where sine is negative), we find sin(α) = -√(187/196) = -√187/14. Next, we need to find the value of cos(β). Since sin(β) = -12/23 and β is in quadrant IV, we can use the Pythagorean identity sin^2(β) + cos^2(β) = 1. Plugging in sin(β) = -12/23, we have (-12/23)^2 + cos^2(β) = 1. Simplifying, we get 144/529 + cos^2(β) = 1. Subtracting 144/529 from both sides, we find cos^2(β) = 385/529. Taking the square root of both sides and remembering that β is in quadrant IV (where cosine is positive), we find cos(β) = √(385/529) = √385/23. Finally, we can use the formula for tan(α + β) = (tan(α) + tan(β))/(1 - tan(α)tan(β)). Since tan(α) = sin(α)/cos(α) and tan(β) = sin(β)/cos(β), we can substitute the known values. We find tan(α + β) = (sin(α)/cos(α) + sin(β)/cos(β))/(1 - sin(α)/cos(α) * sin(β)/cos(β)). Substituting the values we found earlier, we have tan(α + β) = (-√187/14)/(√385/23)/(1 - (-√187/14)(-√385/23)). Multiplying the numerator and the denominator by the reciprocal of the denominator, we have tan(α + β) = (-√187/14) * (23/√385) / (1 - (187/14)*(385/23)). Simplifying the fraction, we have tan(α + β) = (-23√187/(14√385)) / (1 - 385/14) = (-23√187/(14√385)) / (14/14 - 385/14). Simplifying the denominator, we have tan(α + β) = (-23√187/(14√385)) / (-371/14). Dividing by a fraction is equivalent to multiplying by the reciprocal, so we get tan(α + β) = (-23√187/(14√385)) * (-14/371). Simplifying further, we have tan(α + β) = 23√187/(371√385). Therefore, tan(α + β) = 23√187/(371√385).

Manny300303199:

Don't know fr

surjithayer:

\[\alpha ~\in~quadrant ~|||\] \[Hence~\cos \alpha~and~\sin \alpha~are~negative.\] \[\sin \alpha=-\sqrt{1-\cos ^2\alpha}=-\sqrt{1-(\frac{ -3 }{ 14 })^2}\] \[=-\sqrt{1-\frac{ 9 }{ 196}}=-\sqrt{\frac{ 187 }{ 196 }}\] \[\beta ~in~quadrant ~iv \] Therefore cos beta is positive. \[\cos \beta=-\sqrt{1-\sin ^2\beta}=\sqrt{1-(\frac{ -12 }{ 23 })^2}\] \[=\sqrt{1-\frac{ 144 }{529 }}=\sqrt{\frac{ 529-144 }{ 529 }}=\sqrt{\frac{ 385 }{ 529 }}\] \[\tan \alpha=\frac{ \sin \alpha }{ \cos \alpha }=\frac{ \frac{ -\sqrt{187}}{ 14} }{ \frac{ -3 }{ 14 } }=\frac{ \sqrt{187} }{ 3 }\] \[\tan \beta=\frac{ \sin \beta }{ \cos \beta } =\frac{ \frac{ -12}{ 23 } }{ \frac{ \sqrt{385} }{ 23 } }=-\frac{ 12 }{\sqrt{385}}\] \[\tan (\alpha+\beta)=\frac{ \tan \alpha+\tan \beta }{ 1-\tan \alpha \tan \beta }\] \[=\frac{ \frac{ \sqrt{187} }{ 3 }+\frac{ -12 }{ \sqrt{385}} }{ 1-\frac{ \sqrt{187} }{ 3 }\times \frac{ -12 }{ \sqrt{385}} }\] \[=\frac{ \frac{ \sqrt{187} \times \sqrt{385}-36}{ 3\sqrt{385} } }{ 1+\frac{ 12\sqrt{187} }{ 3\sqrt{385} } }\] \[=\frac{ \sqrt{71995}-36 }{ 3\sqrt{385}+12\sqrt{187} }\]